In the Stackelberg model we saw in class there were two firms 1 and 2. Suppose that the market demand is p(Q) = 60−Q, where as in class Q is the aggregate quantity. The const function for firm 1 is c1(q1) = 10q1 and the cost function for firm 2 is c2(q2) = q2. Firm 1 is the leader and Firm 2 is the follower.

(a) Solve for the follow’s reaction function, and the leader’s maximization problem.

(b) Describe the Stackelberg equilibrium and solve for the Stackelberg equilibrium outcome. Compute each firm’s profit.
Answer #1

follower, firm 2, 80 Q) | P(Q) = 60-0, Q = 9+92, MG=10, MG=1. a) To = (P-MC2) 92 T2 = (60-9-92-1) 92 T2 = (59.69 (59-W9292 Fo